OpenStudy (anonymous):

Simplify. -11*4/(-2) and 16/(-2)-3

OpenStudy (anonymous):

okay what is 11*4?

OpenStudy (gw2011):

-11*4/(-2)=(-44)/(-2)=22 16/-(2)-3=-8-3=-11